# Lecture Notes on Spectral Independence and Bases of a Matroid: Local-to-Global and Trickle-Down from a Markov Chain Perspective

### v3

@@ 4 Rapid Mixing from Spectral Independence
The main result of this section is Theorem 1, which states that if $\mu$ is $\eta$-spectrally independent for a constant $\eta$, then the Glauber dynamics has polynomial mixing time.
@@ 4.2 Rapid Mixing Theorem
$\gamma(\mathrm{P_{Gl}})\geq\frac{C(\eta)}{n^{1+\eta}},$ (5)
for a constant $C(\eta)>0$. By Eq. 4 this implies relaxation time $O(n^{1+\eta})$ and mixing time $O(n^{2+\eta}\log{n})$, and hence this proves Theorem 1.
@@ 4.3 Local to Global: Random Walk Theorem of Alev-Lau
The key technical tool in the proof of Eq. 5 is the following local-to-global theorem of Alev and Lau [AL20], which improved upon Kaufman and Oppenheim [KO20]. Our input graph $G=(V,E)$ has $n=|V|$ vertices and each vertex has two possible spin assignments 00 or $1$. The state space of the local walk are the $2n$ (vertex,spin) pairs $(v_{i},s_{i})\in V\times\{0,1\}$. The transition probabilities are defined based on their pairwise joint probabilities. The stationary distribution of the local walk corresponds to the marginal probability $\mu(\sigma(v_{i})=s_{i})$ of that particular vertex $v_{i}$ having the specified spin $s_{i}$.
@@
$Q((i,s_{i}),(j,s_{j}))=\frac{1}{n-1}\Pr_{\sigma\sim\mu}[\sigma(j)=s_{j}\mid\sigma(i)=s_{i}],$
and when $i=j$, then for $s,s^{\prime}\in\{0,1\}$, let $Q((i,s),(i,s^{\prime}))=0$. Moreover, for $20\leq k\leq n-$ 2, for a subset $S\subset V$ where $|S|=k$, for a pinning $\tau:S\rightarrow\{0,1\}$ where $\tau\in\mathcal{P}$, for $i,j\notin S$ where $i\neq j$, and $s_{i},s_{j}\in\{0,1\}$, let
$Q_{\tau}((i,s_{i}),(j,s_{j}))$ $=$ $\frac{1}{n-k-1}\Pr_{\sigma\sim\mu}(\sigma(j)=s_{j}\mid\sigma(i)=s_{i},\sigma(S)=\tau)$ (6)
$=$ $\frac{1}{n-k-1}\Pr_{\sigma\sim\mu_{\tau}}(\sigma(j)=s_{j}\mid\sigma(i)=s_{i}),$
and when $i=j$ we set the entries to 00 as before. Notice that $Q=Q_{\emptyset}$, that is the local walk $Q$ corresponds to the empty pinning. In the definition of the local walk $Q$ and the generalization in Eq. 6, we restrict attention to pairs $(i,s_{i})$ where $\mu_{\tau}(\sigma(i)=s_{i})>0$.
@@
The upcoming Random Walk Theorem of Alev and Lau [AL20] is referred to as a local-to-global theorem because it relates the behavior of local chains $Q$ to the Glauber dynamics which is a global chain. To do this we need to consider the local walk on each level. Namely, for $k$ where $20\leq k\leq n-$ 2, for $\tau\in\mathcal{P}_{k}$, consider the local walk $Q_{\tau}$. Let $\gamma_{k}$ be the minimum spectral gap for a local walk $Q_{\tau}$ where $\tau\in\mathcal{P}_{k}$.
@@
Theorem 4 (Random Walk Theorem [AL20]).
@@ 4.4 Local Walk Connection to Influence Matrix
[…]
@@
Consider the local random walk $Q$. Notice that this walk has an $n$-partite structure, namely, for every $v_{i}\in V$, the states $(v_{i},0)$ and $(v_{i},1)$ are an independent set, and each of these $n$ independent sets are fully connected by a (weighted) bipartite clique. This $n$-partite structure yields eigenvalues $-1/(n-1)$ for $Q$. The stationary distribution for $Q$ is an eigenvector with eigenvalue $1$ (note that when $n\geq 3$ then the chain is clearly aperiodic and hence ergodic).
Let $\pi$ denote the stationary distribution of the local walk $Q$. For a vertex $j\in V$ and a spin/label $s_{j}\in\{0,1\}$, we have $\pi((j,s_{j}))=(1/n)\Pr_{\sigma\sim\mu}[\sigma(j)=s_{j}]$, and hence the distribution $\pi$ corresponds to the normalized marginal distribution of the Gibbs distribution $\mu$. Moreover, let $\pi_{j}$ denote the marginal distribution at vertex $j$; that is, $\pi_{j}((j,s_{j}))=\Pr_{\sigma\sim\mu}[\sigma(j)=s_{j}]$ for $s_{j}\in\{0,1\}$, and the remaining entries set to 00.
We will “zero-out” these trivial eigenvalues $1$ and $-1/(n-1)$ in the following manner:
@@
$M_{\pi}((i,s_{i}),(j,s_{j}))=\frac{1}{n-1}\Big{(}\Pr_{\sigma\sim\mu}[\sigma(j)=s_{j}\mid\sigma(i)=s_{i}]-\Pr_{\sigma\sim\mu}[\sigma(j)=s_{j}]\Big{)},$
@@
Note that if $w$ is a left-eigenvector of $A_{\pi}-B_{\pi}$ then $\bigl{(}\begin{smallmatrix}w\\ -w\end{smallmatrix}\bigr{)}$ is a left-eigenvector of $M_{\pi}$ with the same eigenvalue. Moreover, vectors of the form $(v^{T}v^{T})$ (a space of dimension $n$) are both: right-eigenvectors of $M_{\pi}$ with eigenvalue 00, and are perpendicular to the left-eigenvectors of the form $\bigl{(}\begin{smallmatrix}w\\ -w\end{smallmatrix}\bigr{)}$. These vectors of the form $(v^{T}v^{T})$ yield right-eigenvectors of $\frac{n}{n-1}\mathbf{1}\pi^{T}-\frac{1}{n-1}\sum_{i=1}^{n}\mathbf{1}_{i}\pi_{i}^{T}$, where $\mathbf{1}$ has eigenvalue $1$ and the subspace perpendicular to $\pi$ (a space of dimension $n-1$) has eigenvectors with eigenvalue $-\frac{1}{n-1}$. This implies (9).
This completes the proof of the lemma for the case without pinning. The proof easily generalizes to an arbitrary pinning $\tau\in\mathcal{P}_{k}$ with $(n-1)$ replaced by $(n-k-1)$. ∎
4.5 Proof of Rapid Mixing (Theorem 1)
We can now utilize Lemma 4.1 and Theorem 4 to conclude $\mathrm{poly}(n)$ relaxation time for the Glauber dynamics and thereby prove Theorem 1.
Proof of Theorem 1.
@@ 4.5 Proof of Rapid Mixing (Theorem 1)
[…]
Now we can apply (7) in the Random Walk Theorem (Theorem 4) and we obtain the following:
@@
where $C,C^{\prime},C^{\prime\prime},C^{\prime\prime\prime}$ are positive constants (which may depend on $\eta$). ∎
5 Random Walk Theorem: Proof of Theorem 4
In this section we will prove the Random Walk Theorem (Theorem 4). We will use random walks between adjacent levels of the associated simplicial complex which corresponds to walks between sets $\mathcal{P}_{k}$ and $\mathcal{P}_{k+1}$. These will be denoted as up and down walks, and their composition as up-down and down-up walks.
@@ 5.1 Up and Down Walks
for $\mbox{for }\tau\in\mathcal{P}_{k},\ \ \ \ \pi_{k}(\tau)=\frac{1}{{n\choose k}}\mu(\tau),$
@@
$\sum_{\eta\in\mathcal{P}_{r}:\sigma\subset\eta}\mu(\eta)={n-k\choose r-k}\mu(\sigma).$
@@
$\pi_{k}(\sigma)=\frac{\mu(\sigma)}{{n\choose k}}=\frac{1}{{n-k\choose r-k}}\sum_{\eta\in\mathcal{P}_{r}:\sigma\subset\eta}\frac{\mu(\eta)}{{n\choose k}}=\frac{1}{{r\choose k}}\sum_{\eta\in\mathcal{P}_{r}:\sigma\subset\eta}\frac{\mu(\eta)}{{n\choose r}}={r\choose k}^{-1}\sum_{\eta\in\mathcal{P}_{r}:\sigma\subset\eta}\pi_{r}(\eta).$ (15)
@@
$\mbox{for }\tau\in\mathcal{P}_{\eta,j}\mbox{ where }\tau:S^{\prime}\rightarrow\{0,1\}\mbox{ with }S^{\prime}\subset V\setminus S,\ \ \ \ \pi_{\eta,j}(\tau)=\frac{1}{{n-\ell\choose j}}\mu(\sigma(S^{\prime})=\tau\mid\sigma(S)=\eta).$
@@
$\sum_{\begin{subarray}{c}(j^{\prime},s_{j^{\prime}}):\\ j^{\prime}\notin S,s^{\prime}_{j}\in\{0,1\}\end{subarray}}\pi_{k+1}(\tau\cup(j^{\prime},s_{j^{\prime}}))=\frac{1}{{n\choose k+1}}\sum_{j^{\prime}\notin S}\sum_{s_{j^{\prime}}\in\{0,1\}}\mu(\tau\cup(j^{\prime},s_{j^{\prime}}))\\ =\frac{1}{{n\choose k+1}}\sum_{j^{\prime}\notin S}\mu(\tau)=\frac{n-k}{{n\choose k+1}}\mu(\tau)=(k+1)\pi_{k}(\tau).$
@@
$\pi_{k}(\tau)\mathrm{P}_{k}^{\uparrow}(\tau,\tau\cup\{(j,s_{j}\})=\pi_{k+1}(\tau\cup(j,s_{j}))\mathrm{P}_{k+1}^{\downarrow}.$
@@
Observe that the stationary distribution of $\mathrm{P}_{k}^{\wedge}$ and $\mathrm{P}_{k}^{\vee}$ is $\pi_{k}$.
@@
$=\frac{1}{2}\frac{n}{{n\choose 2}}\times\frac{\mu((i,s_{i})\cup(j,s_{j}))}{2\mu(i,s_{i}))}$
@@ 5.2 Spectrum of Up-Down Chains
$spectrum_{\neq 0}(\mathrm{P}_{k-1}^{\wedge})=spectrum_{\neq 0}(\mathrm{P}_{k}^{\vee}).$
@@
Note that $\mathrm{P}_{k-1}^{\wedge}=P^{\uparrow}_{k-1}P^{\downarrow}_{k}$ and $\mathrm{P}_{k}^{\vee}=P^{\downarrow}_{k}P^{\uparrow}_{k+1}$. Hence, the lemma follows from the general linear algebra fact that for an $n\times m$ matrix $A$ and $m\times n$ matrix $B$ then the non-zero spectrum of $AB$ is the same as the non-zero spectrum of $BA$. For a more detailed proof see [Mou22, Corollary 3.3.2].
@@ 5.3 Proof Setup
We can now restate the Random Walk Theorem in terms of the up-down chains. We will prove, for all $2\leq k\leq n$:
$\gamma(\mathrm{P}_{k}^{\vee})\geq\frac{1}{k}\prod_{i=0}^{k-2}\gamma_{i}.$ (18)
Note, the claim trivially holds for $k=1$ as $\gamma(\mathrm{P}_{1}^{\vee})=1$. Since the Glauber dynamics is equivalent to $\mathrm{P}_{n}^{\vee}$, then the special case of Eq. 7 where $k=n$ is Theorem 4.

### v4

@@ 4 Rapid Mixing from Spectral Independence
The main result of this section is Theorem 1.4, which states that if $\mu$ is $\eta$-spectrally independent for a constant $\eta$, then the Glauber dynamics has polynomial mixing time.
@@ 4.2 Rapid Mixing Theorem
$\gamma(\mathrm{P_{\textsc{gd}}})\geq\frac{C(\eta)}{n^{1+\eta}},$ (31)
for a constant $C(\eta)>0$. By Eq. 16 this implies relaxation time $O(n^{1+\eta})$ and mixing time $O(n^{2+\eta})$, and hence this proves Theorem 1.4.
@@ 4.3 Local to Global: Random Walk Theorem of Alev-Lau
The key technical tool in the proof of Eq. 31 is the following local-to-global theorem of Alev and Lau [AL20], which improved upon Kaufman and Oppenheim [KO20]. Our input graph $G=(V,E)$ has $n=|V|$ vertices and each vertex has two possible spin assignments 0 or $1$. The state space of the local walk are the $2n$ (vertex,spin) pairs $(v_{i},s_{i})\in V\times\{0,1\}$. The transition probabilities are defined based on their pairwise joint probabilities. The stationary distribution of the local walk corresponds to the marginal probability $\mu(\sigma(v_{i})=s_{i})$ of that particular vertex $v_{i}$ having the specified spin $s_{i}$.
@@
$Q((i,s_{i}),(j,s_{j}))=\frac{1}{n-1}\mathrm{Pr}_{\sigma\sim\mu}[\sigma(j)=s_{j}\mid\sigma(i)=s_{i}],$
and when $i=j$, then for $s,s^{\prime}\in\{0,1\}$, let $Q((i,s),(i,s^{\prime}))=0$. Moreover, for $0\leq k\leq n-2$, for a subset $S\subset V$ where $|S|=k$, for a pinning $\tau:S\rightarrow\{0,1\}$ where $\tau\in\mathcal{P}$, for $i,j\notin S$ where $i\neq j$, and $s_{i},s_{j}\in\{0,1\}$, let
$Q_{\tau}((i,s_{i}),(j,s_{j}))$ $=$ $\frac{1}{n-k-1}\mathrm{Pr}_{\sigma\sim\mu}(\sigma(j)=s_{j}\mid\sigma(i)=s_{i},\sigma(S)=\tau)$ (32)
$=$ $\frac{1}{n-k-1}\mathrm{Pr}_{\sigma\sim\mu_{\tau}}(\sigma(j)=s_{j}\mid\sigma(i)=s_{i}),$
and when $i=j$ we set the entries to 0 as before. Notice that $Q=Q_{\emptyset}$, that is the local walk $Q$ corresponds to the empty pinning. In the definition of the local walk $Q$ and the generalization in Eq. 32, we restrict attention to pairs $(i,s_{i})$ where $\mu_{\tau}(\sigma(i)=s_{i})>0$.
@@
The upcoming Random Walk Theorem of Alev and Lau [AL20] is referred to as a local-to-global theorem because it relates the behavior of local chains $Q$ to the Glauber dynamics which is a global chain. To do this we need to consider the local walk on each level. Namely, for $k$ where $0\leq k\leq n-2$, for $\tau\in\mathcal{P}_{k}$, consider the local walk $Q_{\tau}$. Let $\gamma_{k}$ be the minimum spectral gap for a local walk $Q_{\tau}$ where $\tau\in\mathcal{P}_{k}$.
@@
Theorem 4.1 (Random Walk Theorem [AL20]).
@@ 4.4 Local Walk Connection to Influence Matrix
$\mbox{spectrum}(Q_{\tau})=\mbox{spectrum}\left(\frac{1}{n-k-1}\left(\Psi_{\tau}-I\right)\right)\cup\{1\}\cup\left\{n-k-1\mbox{ copies of }\frac{-1}{n-k-1}\right\}.$ (35)
@@
Consider the local random walk $Q$. Notice that this walk has an $n$-partite structure, namely, for every $v_{i}\in V$, the states $(v_{i},0)$ and $(v_{i},1)$ are an independent set, and each of these $n$ independent sets are fully connected by a (weighted) bipartite clique. This $n$-partite structure yields eigenvalues 1/(n1)-1/(n-1) for $Q$. The stationary distribution for $Q$ is an eigenvector with eigenvalue $1$ (note that when $n\geq 3$ then the chain is clearly aperiodic and hence ergodic).
Let $\pi$ denote the stationary distribution of the local walk $Q$. For a vertex $j\in V$ and a spin/label $s_{j}\in\{0,1\}$, we have $\pi((j,s_{j}))=(1/n)\mathrm{Pr}_{\sigma\sim\mu}[\sigma(j)=s_{j}]$, and hence the distribution $\pi$ corresponds to the normalized marginal distribution of the Gibbs distribution $\mu$. Moreover, let $\pi_{j}$ denote the marginal distribution at vertex $j$; that is, $\pi_{j}((j,s_{j}))=\mathrm{Pr}_{\sigma\sim\mu}[\sigma(j)=s_{j}]$ for $s_{j}\in\{0,1\}$, and the remaining entries set to 0.
We will “zero-out” these trivial eigenvalues $1$ and 1/(n1)-1/(n-1) in the following manner:
@@
$M_{\pi}((i,s_{i}),(j,s_{j}))=\frac{1}{n-1}\Big{(}\mathrm{Pr}_{\sigma\sim\mu}[\sigma(j)=s_{j}\mid\sigma(i)=s_{i}]-\mathrm{Pr}_{\sigma\sim\mu}[\sigma(j)=s_{j}]\Big{)},$
@@
Note that if $w$ is a left-eigenvector of $A_{\pi}-B_{\pi}$ then $\bigl{(}\begin{smallmatrix}w\\ -w\end{smallmatrix}\bigr{)}$ is a left-eigenvector of $M_{\pi}$ with the same eigenvalue. Moreover, vectors of the form $(v^{T}v^{T})$ (a space of dimension $n$) are both: right-eigenvectors of $M_{\pi}$ with eigenvalue 0, and are perpendicular to the left-eigenvectors of the form $\bigl{(}\begin{smallmatrix}w\\ -w\end{smallmatrix}\bigr{)}$. These vectors of the form $(v^{T}v^{T})$ yield right-eigenvectors of $\frac{n}{n-1}\mathbf{1}\pi^{T}-\frac{1}{n-1}\sum_{i=1}^{n}\mathbf{1}_{i}\pi_{i}^{T}$, where $\mathbf{1}$ has eigenvalue $1$ and the subspace perpendicular to $\pi$ (a space of dimension n1n-1) has eigenvectors with eigenvalue $\frac{1}{n-1}$. This implies (35).
This completes the proof of the lemma for the case without pinning. The proof easily generalizes to an arbitrary pinning $\tau\in\mathcal{P}_{k}$ with (n1)(n-1) replaced by (nk1)(n-k-1). ∎
4.5 Proof of Rapid Mixing (Theorem 1.4)
We can now utilize Lemma 4.2 and Theorem 4.1 to conclude $\mathrm{poly}(n)$ relaxation time for the Glauber dynamics and thereby prove Theorem 1.4.
Proof of Theorem 1.4.
@@
 $\mbox{ if }k\geq n-\eta-1,\mbox{ then }\gamma_{k}\geq C\mbox{ for some constant $C>0$.}$ (39)
Now we can apply (33) in the Random Walk Theorem (Theorem 4.1) and we obtain the following:
@@
where $C,C^{\prime},C^{\prime\prime},C^{\prime\prime\prime}$ are positive constants (which may depend on $\eta$).
We have shown that the spectral gap of the Glauber dynamics is $\Omega(1/n^{1+\eta})$. Applying Eq. 16 we obtain that the relaxation time satisfies $T_{\mathrm{relax}}=O(n^{1+\eta})$. Moreover, since $\log(\mu^{*})\leq O(n\log{q})$ and we are considering the case $q=2$, then, from Eq. 16, we have the mixing time bound of $T_{\mathrm{mix}}=O(n^{2+\eta})$. This establishes the upper bounds on the relaxation time and mixing time stated in Theorem 1.4. ∎
5 Random Walk Theorem: Proof of Local-to-Global
In this section we will prove the Random Walk Theorem (Theorem 4.1). We will use random walks between adjacent levels of the associated simplicial complex which corresponds to walks between sets $\mathcal{P}_{k}$ and $\mathcal{P}_{k+1}$. These will be denoted as up and down walks, and their composition as up-down and down-up walks.
@@ 5.1 Up and Down Walks
$\mbox{for }\tau\in\mathcal{P}_{k},\ \ \ \ \pi_{k}(\tau)=\frac{1}{\binom{n}{k}}\mu(\tau),$
@@
$\sum_{\eta\in\mathcal{P}_{r}:\sigma\subset\eta}\mu(\eta)=\binom{n-k}{r-k}\mu(\sigma).$
@@
$\pi_{k}(\sigma)=\frac{\mu(\sigma)}{{\binom{n}{k}}}=\frac{1}{\binom{n-k}{r-k}}\sum_{\eta\in\mathcal{P}_{r}:\sigma\subset\eta}\frac{\mu(\eta)}{{\binom{n}{k}}}=\frac{1}{\binom{r}{k}}\sum_{\eta\in\mathcal{P}_{r}:\sigma\subset\eta}\frac{\mu(\eta)}{\binom{n}{r}}={\binom{r}{k}}^{-1}\sum_{\eta\in\mathcal{P}_{r}:\sigma\subset\eta}\pi_{r}(\eta).$ (41)
@@
$\mbox{for }\tau\in\mathcal{P}_{\eta,j}\mbox{ where }\tau:S^{\prime}\rightarrow\{0,1\}\mbox{ with }S^{\prime}\subset V\setminus S,\ \ \ \ \pi_{\eta,j}(\tau)=\frac{1}{\binom{n-\ell}{j}}\mu(\sigma(S^{\prime})=\tau\mid\sigma(S)=\eta).$
@@
$\sum_{\begin{subarray}{c}(j^{\prime},s_{j^{\prime}}):\\ j^{\prime}\notin S,s^{\prime}_{j}\in\{0,1\}\end{subarray}}\pi_{k+1}(\tau\cup(j^{\prime},s_{j^{\prime}}))=\frac{1}{\binom{n}{k+1}}\sum_{j^{\prime}\notin S}\sum_{s_{j^{\prime}}\in\{0,1\}}\mu(\tau\cup(j^{\prime},s_{j^{\prime}}))\\ =\frac{1}{\binom{n}{k+1}}\sum_{j^{\prime}\notin S}\mu(\tau)=\frac{n-k}{\binom{n}{k+1}}\mu(\tau)=(k+1)\pi_{k}(\tau).$
@@
$\pi_{k}(\tau)\mathrm{P}_{k}^{\uparrow}(\tau,\tau\cup(j,s_{j}))=\pi_{k+1}(\tau\cup(j,s_{j}))\mathrm{P}_{k+1}^{\downarrow}(\tau\cup(j,s_{j}),\tau).$ (42)
@@
Observe that the stationary distribution of $\mathrm{P}_{k}^{\wedge}$ and $\mathrm{P}_{k}^{\vee}$ is $\pi_{k}$. The reversibility condition Eq. 42 shows that the two operators $\mathrm{P}_{k}^{\downarrow},\mathrm{P}_{k-1}^{\uparrow}$ are adjoint to each other, which implies that both up-down walks and down-up walks are reversible and positive semidefinite (PSD); see [CCGP25, Section 2] for more discussion.
@@
$=\frac{1}{2}\frac{n}{\binom{n}{2}}\times\frac{\mu((i,s_{i})\cup(j,s_{j}))}{2\mu(i,s_{i}))}$
@@ 5.2 Spectrum of Up-Down Chains
$\mathrm{spectrum}_{\neq 0}(\mathrm{P}_{k-1}^{\wedge})=\mathrm{spectrum}_{\neq 0}(\mathrm{P}_{k}^{\vee}).$
@@
Note that $\mathrm{P}_{k-1}^{\wedge}=P^{\uparrow}_{k-1}P^{\downarrow}_{k}$ and $\mathrm{P}_{k}^{\vee}=P^{\downarrow}_{k}P^{\uparrow}_{k+1}$. Hence, the lemma follows from the general linear algebra fact that for an $n\times m$ matrix $A$ and $m\times n$ matrix $B$ then the non-zero spectrum of $AB$ is the same as the non-zero spectrum of $BA$. For a more detailed proof see [Mou22, Corollary 3.3.2]. ∎
@@ 5.3 Proof Setup
We can now state a more general version of the Random Walk Theorem (Theorem 4.1) which bounds the spectral gap of the down-up chain at any level. We will prove, for all $2\leq k\leq n$:
$\gamma(\mathrm{P}_{k}^{\vee})\geq\frac{1}{k}\prod_{i=0}^{k-2}\gamma_{i},$ (45)
where
$\gamma_{i}=\min_{\tau\in\mathcal{P}_{i}}\gamma(Q_{\tau})$
is the spectral gap for the local walk with a worst-case pinning of $i$ vertices. Note, the claim trivially holds for $k=1$ as $\gamma(\mathrm{P}_{1}^{\vee})=1$. Since the Glauber dynamics is equivalent to $\mathrm{P}_{n}^{\vee}$, then Theorem 4.1 is the special case of Eq. 33 where $k=n$.